If 88 liters hydrogen gas reacts how many grams of aluminum will be produced


1
Expert's answer
2021-03-29T06:02:28-0400

2AlCl3 + 3H2 -> 2Al + 6HCl

2 m of Al = 3 m of H2

Molar Mass of Al = 26.981539

Molar Mass of H2 = 2.016

= 88000/2.016 = 43650.794 moles of H2

2/3×43650.794 = 29100.53 moles of Al

= 29100.53 ×26.981539

= 785177.067 g of aluminum